OpenStudy (anonymous):

Sin x cos x- sin^2 x= 0

OpenStudy (anonymous):

57.3 plusorminus 2 pi n

OpenStudy (anonymous):

How do you solve by factoring?

OpenStudy (vishweshshrimali5):

\[\large{\sin x \cos x - \sin^2 x = 0}\] \[\large{\implies \sin x(\cos x - \sin x) = 0}\] Now: \[\large{a\times b = 0}\] \[\large{\implies a = 0 \ or \ b = 0}\]

OpenStudy (vishweshshrimali5):

Get it ?

OpenStudy (anonymous):

I'm still a little confuse how you got that

OpenStudy (anonymous):

\[ \sin x = 0 \quad \text{or}\quad \cos x - \sin x = 0 \]

OpenStudy (anonymous):

\[ \color{red}{\sin x} \color{blue}{\cos x} - \color{purple}{\sin^2 x }= \color{red}{\sin x} \color{blue}{\cos x} - \color{red}{\sin x} \color{blue}{\sin x} =\color{red}{ \sin x} (\color{blue}{\cos x - \sin x}) \]

OpenStudy (anonymous):

So would that be the answer?